OpenStudy (anonymous):

@Mathematics For spring break this year your family decided to visit Michigan, but it snows the first weekend you are there. You decide to make a snowman. If the bottom ball is 3 feet across, the middle ball is 2 feet across, and the top is 1 foot across, how much snow is needed to build the whole snowman? Round your answer to the nearest cubic foot. A) 19 cubic feet of snow B) 31 cubic feet of snow C) 511 cubic feet of snow D) 1022 cubic feet of snow

OpenStudy (mathteacher1729):

Across = "The diameter", right? if yes, then the volume of a sphere with diameter d is \[v = \frac{4}{3}\left(\frac{d}{2}\right)^3\]

OpenStudy (mathteacher1729):

So find the volume of each of the three parts of the snowman, and add them together to find the total volume. :)

OpenStudy (anonymous):

I think it's A

OpenStudy (anonymous):

@mathteacher1729 what do you think the answer choice is ?

OpenStudy (mathteacher1729):

I know what the correct answer is. I'm trying to help you solve the problem yourself though. :) Do you know how to substitute the values d = 1, 2, and 3 into the formula for the volume of each sphere?

OpenStudy (mathteacher1729):

\[v(1) = \frac{4}{3}\left(\frac{1}{2}\right)^3 \approx\] \[v(2) = \frac{4}{3}\left(\frac{2}{2}\right)^3 \approx\] \[v(3) = \frac{4}{3}\left(\frac{3}{2}\right)^3 \approx\] Now add up \(v(1) + v(2) + v(3)\) "volume of sphere with diameter 1 , volume of sphere with diameter 2 , volume of sphere with diameter 3" :)

OpenStudy (anonymous):

you have lost \(\pi\)

OpenStudy (anonymous):

I understand what you're trying to do,, but I'm not sure how to do it or what you're talking about far as solving the problem with the equations you have set up

OpenStudy (anonymous):

Yes, you left the \[\pi\]

OpenStudy (anonymous):

formula is \[V=\frac{4}{3}\pi R^3=\frac{4}{3}\pi \left(\frac{d}{2}\right)^3\]